OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (anonymous):

hey b-tich

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

hey Krissie can you help me please

OpenStudy (welshfella):

one zero is complex 1 + 3i so another zero will be its conjugate do you know what that is?

OpenStudy (anonymous):

um hello? and no i don't

OpenStudy (welshfella):

the 2 values stay the same but the sign changes so the conjugate is 1 - 3i

OpenStudy (welshfella):

you first write the polynomial in factored form one of the roots is 2 so what would one factor be?

OpenStudy (welshfella):

recall quadratic equations (x - 1)(x + 2) = 0 what are the zeros for this equation?

OpenStudy (welshfella):

have you done these before?

OpenStudy (anonymous):

nope, these are my answers and i think its.. the first one f(x) = x4 - 2x2 + 36x - 80 f(x) = x4 - 3x3 + 6x2 - 18x + 80 f(x) = x4 - 9x2 + 36x - 80 f(x) = x4 - 3x3 - 6x2 + 18x - 80

OpenStudy (welshfella):

if you haven't done quadratics I'm surprised you are given this problem you could try substituting each zero - that is 2, -4 , 1 + 3i and 1 - 3i into each option - and see which one they fit

OpenStudy (anonymous):

can you please just tell me which one it is?